# PartNeXt: A Next-Generation Dataset for Fine-Grained and Hierarchical 3D Part Understanding

### v1

@@ 6 Acknowledgement
This work was sponsored by Shanghai Pujiang Program (24PJA080). We also appreciate Benyuan AI Data for providing support in data annotation, as well as the 35 annotators for their valuable contributions.
@@ Appendix C Comparison with Recent 3D Part Dataset
We present a comparison between PartNeXt and recent 3D part-level datasets, as shown in Table 7. This includes detailed information on several 3D object part-level datasets such as PartNet[32] and PartObjaverseTiny from SAMPart3D[48].
@@
The PartNet[32] dataset is annotated directly on textureless models and requires remeshing during the annotation process. As a result, the final annotations lack both texture and original geometry, omitting important color information and potential geometric details (such as mesh topology and face distribution), which are crucial for comprehensive 3D understanding. Besides, PartNet requires manually drawn cutting lines after remeshing to achieve segmentation. As a result, the boundaries of the parts are often not smooth. In contrast, our method leverages per-face annotations, effectively avoiding these issues, the comparison is shown in Figure 6.
@@
PartObjaverseTiny[48] retains the original geometry during annotation, but it lacks a hierarchical structure, making it unsuitable for models that require multi-granularity segmentation (e.g., PointSAM[53]).
In terms of scale, our dataset is comparable to PartNet[32] in quantity, whereas PartObjaverseTiny[48] contains only 200 models—insufficient for large-scale 3D applications. As for category diversity, our dataset includes more than twice the number of categories compared to PartNet[32]. Although PartObjaverseTiny[48] claims to have subcategories under its eight main categories, the limited quantity constrains its scalability for broader 3D model tasks.
@@ Appendix D Detail of Segmentation BenchMark
We evaluate each method using its publicly available codebase, and follow their original settings and hypermeters. SAMesh [39] utilizes textureless meshes, PartField [25] uses colorless point clouds, sampled from mesh surfaces, SAMPart3D [48] requires point clouds with color and normal attributes, so we sample colorized point clouds from textured meshes as input. Following the evaluation code of PartField, we convert the predicted masks from all methods into a unified format before computing the mIoU.
@@ Appendix E Detail of LLM BenchMark
In all experiments, we use publicly available code. However, we observe that current LLM for 3D understanding shows limited capabilities in instruction following and response formatting. Even with carefully designed prompts, these models struggle to generate well-formatted outputs. To address this issue, we utilize a language model for output postprocess. We deploy the Qwen3-14B[47] AWQ quantized model using the vLLM framework[16] on 4 NVIDIA GeForce 2080Ti GPUs for postprocess.
@@ E.3 Detail of Part Grounding Task
For ShapeLLM[35], we follow the original evaluation code, requiring the model to output the coordinates of the eight vertices of the bounding box. We then compute the minimum and maximum values to calculate the IoU. An example of the output from ShapeLLM[35], along with the ground truth, is shown in Figure  8. These results indicate that current 3D language models are still in the early stages when it comes to the Part Grounding task.
@@
In Figure 9, we present several output examples from 3DLLM[11] and PointLLM[46]. As shown, both types of models struggle to generate precise bounding box coordinates, indicates the weak ability of part grounding.

### v2

@@ 6 Acknowledgement
This work was sponsored by Shanghai Pujiang Program (24PJA080), and supported by HPC Platform of ShanghaiTech University. We also appreciate Benyuan AI Data for providing support in data annotation, as well as the 35 annotators for their valuable contributions.
@@ Appendix C Comparison with Recent 3D Part Dataset
We present a comparison between PartNeXt and recent 3D part-level datasets, as shown in Table 7. This includes detailed information on several 3D object part-level datasets such as PartNet[mo2019partnet] and PartObjaverseTiny from SAMPart3D[yang2024sampart3d].
@@
The PartNet[mo2019partnet] dataset is annotated directly on textureless models and requires remeshing during the annotation process. As a result, the final annotations lack both texture and original geometry, omitting important color information and potential geometric details (such as mesh topology and face distribution), which are crucial for comprehensive 3D understanding. Besides, PartNet requires manually drawn cutting lines after remeshing to achieve segmentation. As a result, the boundaries of the parts are often not smooth. In contrast, our method leverages per-face annotations, effectively avoiding these issues, the comparison is shown in Figure 6.
@@
PartObjaverseTiny[yang2024sampart3d] retains the original geometry during annotation, but it lacks a hierarchical structure, making it unsuitable for models that require multi-granularity segmentation (e.g., PointSAM[zhou2024point]).
In terms of scale, our dataset is comparable to PartNet[mo2019partnet] in quantity, whereas PartObjaverseTiny[yang2024sampart3d] contains only 200 models—insufficient for large-scale 3D applications. As for category diversity, our dataset includes more than twice the number of categories compared to PartNet[mo2019partnet]. Although PartObjaverseTiny[yang2024sampart3d] claims to have subcategories under its eight main categories, the limited quantity constrains its scalability for broader 3D model tasks.
@@ Appendix D Detail of Segmentation BenchMark
We evaluate each method using its publicly available codebase, and follow their original settings and hypermeters. SAMesh [tang2024segment] utilizes textureless meshes, PartField [liu2025partfield] uses colorless point clouds, sampled from mesh surfaces, SAMPart3D [yang2024sampart3d] requires point clouds with color and normal attributes, so we sample colorized point clouds from textured meshes as input. Following the evaluation code of PartField, we convert the predicted masks from all methods into a unified format before computing the mIoU.
@@ Appendix E Detail of LLM BenchMark
In all experiments, we use publicly available code. However, we observe that current LLM for 3D understanding shows limited capabilities in instruction following and response formatting. Even with carefully designed prompts, these models struggle to generate well-formatted outputs. To address this issue, we utilize a language model for output postprocess. We deploy the Qwen3-14B[yang2025qwen3] AWQ quantized model using the vLLM framework[kwon2023efficient] on 4 NVIDIA GeForce 2080Ti GPUs for postprocess.
@@ E.3 Detail of Part Grounding Task
For ShapeLLM[qi2024shapellm], we follow the original evaluation code, requiring the model to output the coordinates of the eight vertices of the bounding box. We then compute the minimum and maximum values to calculate the IoU. An example of the output from ShapeLLM[qi2024shapellm], along with the ground truth, is shown in Figure  8. These results indicate that current 3D language models are still in the early stages when it comes to the Part Grounding task.
@@
In Figure 9, we present several output examples from 3DLLM[hong20233d] and PointLLM[xu2024pointllm]. As shown, both types of models struggle to generate precise bounding box coordinates, indicates the weak ability of part grounding.